# Entering the overcritical regime of nonlinear Breit-Wheeler pair production in collisions of bremsstrahlung $\gamma$-rays and superintense, tightly focused laser pulses

### v1

@@ Abstract
Near-future high-intensity lasers offer prospects for the observation of nonlinear Breit-Wheeler pair production in an overcritical field regime, where the quantum nonlinearity parameter substantially exceeds unity. This experimentally yet unexplored scenario is envisaged here to be reached via the collision of a tightly focused laser pulse with high-energy bremsstrahlung photons. We calculate the achievable number of positrons in a range of laser intensities around 1023 W/cm2 and GeV-energies of the incident bremsstrahlung-generating electron beam. We investigate under which conditions the attenuation of the $\gamma$-beam due to the production process must be taken into account. In the considered interaction regime, where the local production rate grows rather moderately with higher field intensities, it is shown that the range of mostly contributing bremsstrahlung frequencies is generally very broad. For sufficiently large values of the quantum nonlinearity parameter, an optimum domain of frequencies emerges which is located far below the spectral endpoint. Furthermore, we show that it is beneficial for achieving the optimum pair yield to increase the interaction volume by a wider laser focus at the expense of decreased field intensity.
@@ I Introduction
In particular, a nonlinear version of the Breit-Wheeler effect was studied (see [8, 9, 10] and references therein) where pairs are produced by multiphoton absorption in the collision of a high-energy gamma photon with an intense laser wave, according to the reaction $\omega^{\prime}+n\omega\to e^{+}e^{-}$, where $n$ denotes the number of absorbed low-frequency laser photons. Various interaction regimes of the process have been identified that are distinguished by the values of two dimensionless parameters: on the one hand, the classical intensity parameter $\xi=|e\mathcal{E}_{0}|/(m\omega)$, with the laser frequency $\omega$ and its peak field strength $\mathcal{E}_{0}$; and on the other hand, the quantum nonlinearity parameter given by $\kappa=2\omega^{\prime}\mathcal{E}_{0}/(mE_{c})$, with the critical Schwinger field $E_{c}=m^{2}/|e|\approx 1.3\times 10^{16}\;\rm Vcm^{-1}$ and assuming a counterpropagating beam geometry. Besides, $e$ and $m$ denote the electron charge and mass, respectively. The theoretical treatment in the seminal papers [3, 4, 2] relied on monoenergetic gamma photons and a plane-wave description of the laser field. In the 1990s, the nonlinear Breit-Wheeler process has been observed experimentally in a few-photon regime ($n\sim 5$) at $\xi\lesssim 1$, where the process rate was found to follow a power law of the form $R\sim\xi^{2n}$ [11].
@@
A promising possibility to generate high-energy gamma rays is offered by bremsstrahlung of highly relativistic electrons from high-$Z$ targets, as proposed by Reiss in the 1970s [5], and recently further investigated by several groups [35, 36, 32, 37, 33, 34, 38, 39]. In this situation, the gamma photons are not monoenergetic but span a broad range of energies. This route is followed in the experimental designs for nonlinear Breit-Wheeler pair creation at CALA [13] and DESY [14]. If successful, they could be considered as ’discovery experiments’ for the highly nonlinear regime of the strong field Breit-Wheeler process, which will be followed afterwards by ’precision experiments’ that aim at covering large regions of the parameter space with enhanced accuracy. A detailed theoretical analysis of the experiment planned at CALA has recently been carried out [33, 34], including laser focusing effects and the broad frequency spectrum of bremsstrahlung. This experiment is going to operate in the $\xi\gg 1$ regime with $\kappa\sim 1$, exploiting laser wakefield accelerated incident electrons of about 2.5 GeV energy to generate the bremsstrahlung in a tungsten target [13].
@@
In the present paper, we extend the previous analysis [33] to the overcritical regime of large $\kappa$ values up to $\kappa\approx 30$. It could, for instance, be realized in the future at the Extreme Light Infrastructure (ELI) [15] or CoReLS [40] facilities. Our main goal is to reveal characteristic differences between the intermediate ($\kappa\approx 1$) and overcritical ($\kappa\gg 1$) regimes of the nonlinear strong-field Breit-Wheeler process with focused laser pulses and gamma photons from bremsstrahlung.  [^1] The $\kappa$ values considered here remain far below the domain where perturbation theory with respect to the quantized radiation field is expected to break down, i.e. $\kappa\ll\alpha^{-3/2}\approx 1.6\times 10^{3}$, with $\alpha$ referring to the fine-structure constant [8, 9, 10, 41]. We will show that the overcritical field regime requires an adjusted theoretical approach and exhibits qualitatively distinct properties in terms of its scaling behaviours. This affects, in particular, its dependencies on the laser pulse duration and focusing as well as the mainly contributing spectral range of gamma photons. Our outcomes can be relevant for the design of future experiments.
@@
It is worth noting that the overcritical regime of strong-field pair production can also be probed when an incident high-energy electron beam interacts directly with an ultrastrong laser pulse [12, 42, 43]. In this case, however, the electron beam can be strongly distorted by radiation reaction effects before reaching the focal region of highest field strength. In this regard, the use of a target foil that first converts the incident electron energy in a controled way into $\gamma$ photons, which afterwards penetrate into the laser focal region, offers certain advantages. Figure 1 shows a scheme of the envisaged setup.
@@
Figure 1: Scheme of the experimental setup for entering the overcritical regime of the nonlinear Breit-Wheeler pair production process from the collision of bremsstrahlung gamma photons and a high-intensity laser pulse. The collision of a monoenergetic electron beam with a narrow high-$Z$ target will result in the bremsstrahlung burst utilized in the setup. A magnet after the target will deflect residual electrons away from the site where the photon-photon collision occurs.
@@
Our study is presented as follows: In Sec. II we first briefly review the theoretical description of Breit-Wheeler pair creation in collisions of bremsstrahlung $\gamma$-rays with a focussed high-intensity laser pulse from Ref. [33] and, afterwards, extend it to the overcritical regime. Next, in Sec. III, our numerical results obtained for the overcritical regime $\kappa>1$ are discussed and put in contrast to the properties of the pair creation process at $\kappa\approx 1$. Finally, the conclusions are drawn in the last section. Throughout the manuscript, the Lorentz-Heaviside units [$c=\hbar=\epsilon_{0}=1$], and the metric tensor with the signature $\mathrm{diag}(g^{\mu\nu})=(1,-1,-1,-1)$ are used.
@@ II Theoretical treatment
Equation (1) forms the basis for two different theoretical approaches to strong-field Breit-Wheeler pair production that will be outlined in Secs. II.2 and II.3 below. Beforehand, we shall describe the colliding high-intensity focussed laser pulse and high-energy bremsstrahlung $\gamma$-rays.
II.1 Descriptions of the focussed laser pulse and bremsstrahlung spectrum
@@
where $\tau$ stands for the pulse duration (at FWHM from the intensity), $\rho^{2}=x^{2}+y^{2}$, $\zeta=z/z_{R}$ with the Rayleigh length $z_{R}=\pi w_{0}^{2}/\lambda$ and the beam waist radius at the focus $w_{0}$, whereas $w(z)=w_{0}\sqrt{1+\zeta(z)^{2}}$. Moreover, $E_{x}=B_{y}$ holds and the field phase reads
@@
for a counterpropagating geometry, in which case, the interacting gamma quantum is characterized by a wave vector $k^{\prime}=\omega^{\prime}(1,0,0,-1)$.
In order to take into account bremsstrahlung, the corresponding photon distribution has to be considered. For high energies $E_{0}$ of incident bremsstrahlung electrons—which is the case in the present paper as we assume GeV energies—the collimation angle of the emitted photons can be approximated as an inverse Lorentz factor $\theta_{\gamma}\approx 1/\gamma_{e}=m/E_{0}\sim\mathcal{O}(1)$ mrad. Hence, the vast majority of the produced bremsstrahlung photons is emitted in the direction of propagation of incident electrons. Consequently, the photons energy spectrum can be approximated by [45, 46]
@@
with the normalized photon energy $f=\omega^{\prime}/E_{0}$ and the normalized target thickness $\ell=L_{\mathrm{T}}/L_{\mathrm{rad}}$, which relates the target thickness $L_{\mathrm{T}}$ to the radiation length of the material $L_{\rm rad}$. Note that the formula above is valid within the complete screening approximation and for material with high Z-numbers, as well as for very thin targets with $\ell\ll 1$. Hereafter, a target made up of tungsten is considered with $L_{\rm rad}=3.5$ mm and $L_{T}=50\>\rm\mu m$, corresponding to $\ell=0.015$ [13]. In Fig. 2 one sees the bremsstrahlung spectrum per radiating incident electron as given in Eq. (5) for the chosen target thickness as a function of $f$. We shall assume throughout that the incident electron bunch contains a total (absolute) charge of $Q_{e}=10$ pC and that $\delta=1$% of the electrons will emit bremsstrahlung, which represents a reasonable fraction for the chosen target thickness [13]. Accordingly, the total number of radiating electrons is $Q_{e}\delta/|e|\approx 6\times 10^{5}$.
@@
where $\theta_{\mathrm{rms}}=(\theta_{e^{-}}^{2}+\theta_{\gamma}^{2})^{1/2}$. Here, it is assumed that the electrons are accelerated via laser wakefield acceleration by a portion of the same laser that is used for the pair production. Thus, the longitudinal extension of the electron beam is set to $2z_{R}$. For further details, we refer to Sec. II.1 in [33]. In the further course of our consideration the electron beam spreading parameters are set to $L_{e^{-}}=10$ cm and $\theta_{e^{-}}=0.5$ mrad. The distance to the interaction zone travelled by the $\gamma$-photons is taken as $L=0.5$ m. Accordingly, when the incident electron energy is $E_{0}=2.5$ GeV, the cross sectional area amounts to $A_{\gamma}\approx\pi(320\mu{\rm m})^{2}$. It slightly shrinks to $A_{\gamma}\approx\pi(300\,\mu{\rm m})^{2}$ for $E_{0}=10$ GeV, because the beam of $\gamma$-photons is less divergent then.
@@ II.2 Pair production at moderate field strengths: Stable $\gamma$-beam model
The boundaries of the $z$-integration take the spatial overlap between laser pulse and bremsstrahlung beam into account (see also Fig. 7 below). In order to include the energy distribution of the bremsstrahlung, an incoherent average of Eq. (8) over the spectrum (5) must be taken. Thus, the number of created Breit-Wheeler pairs per radiating electron in our S-model is given by
@@ II.3 Pair production at very high field strengths: Decaying $\gamma$-beam model
The method described in Sec. II.2 is only applicable when the quantum nonlinearity parameter $\kappa$ and the laser pulse duration $\tau$ are not too large. Otherwise this method will overestimate the pair production yield, because it disregards the attenuation of the $\gamma$-rays that arises from their decay into electron positron pairs. In the following we present another calculational method that is able to describe the nonlinear Breit-Wheeler process at $\kappa$ values substantially exceeding unity. To this end, we adjust the treatment put forward in Ref. [32] and extend it to incorporate $\gamma$-rays from bremsstrahlung with a broad frequency spectrum.
@@
which coincides with the probability for producing a pair. In the exponent of Eq. (12), the pair production rate $R(\kappa)$ enters, which is evaluated along the trajectory of the $\gamma$-photon. Clearly, when the time-integrated rate is much smaller than unity, the exponential can be expanded and one obtains
@@
Afterwards, we obtain the total number of produced pairs per radiation electron in the D-model by weighting with the bremsstrahlung spectrum, according to
@@
For our numerical calculations we cut off the integration domain in Eq. (14) at a sufficiently large upper boundary $b_{\rm max}=3w_{0}$, that corresponds to a cross sectional area of $A_{\rm max}=\pi(3w_{0})^{2}$. Then we distribute the associated portion of $A_{\rm max}/A_{\gamma}$ bremsstrahlung $\gamma$-photons uniformly over the area $A_{\rm max}$ and calculate the contribution of each trajectory to the pair yield.

### v2

@@ Abstract
Near-future high-intensity lasers offer prospects for the observation of nonlinear Breit-Wheeler pair production in an overcritical field regime, where the quantum nonlinearity parameter substantially exceeds unity. This experimentally yet unexplored scenario is envisaged here to be reached via the collision of a tightly focused laser pulse with high-energy bremsstrahlung photons. We calculate the achievable number of pairs in a range of laser intensities around 1023 W/cm2 and GeV-energies of the incident bremsstrahlung-generating electron beam. We investigate under which conditions the attenuation of the $\gamma$-beam due to the production process must be taken into account and how much the second generation of created pairs contributes to the total yield. In the considered interaction regime, where the local production rate grows rather moderately with higher field intensities, it is shown that the range of mostly contributing bremsstrahlung frequencies is generally very broad. For sufficiently large values of the quantum nonlinearity parameter, an optimum domain of frequencies emerges which is located far below the spectral endpoint. Furthermore, we show that it is beneficial for achieving the optimum pair yield to increase the interaction volume by a wider laser focus at the expense of decreased field intensity.
@@ I Introduction
In particular, a nonlinear version of the Breit-Wheeler effect was studied (see [8, 9, 10] and references therein) where pairs are produced by multiphoton absorption in the collision of a high-energy $\gamma$-photon with an intense laser wave, according to the reaction $\omega^{\prime}+n\omega\to e^{+}e^{-}$, where $n$ denotes the number of absorbed low-frequency laser photons. Various interaction regimes of the process have been identified that are distinguished by the values of two dimensionless parameters: on the one hand, the classical intensity parameter $\xi=|e\mathcal{E}_{0}|/(m\omega)$, with the laser frequency $\omega$ and its peak field strength $\mathcal{E}_{0}$; and on the other hand, the quantum nonlinearity parameter given by $\kappa=2\omega^{\prime}\mathcal{E}_{0}/(mE_{c})$, with the critical Schwinger field $E_{c}=m^{2}/|e|\approx 1.3\times 10^{16}$ V/cm and assuming a counterpropagating beam geometry. Besides, $e$ and $m$ denote the electron charge and mass, respectively. The theoretical treatment in the seminal papers [3, 4, 2] relied on monoenergetic $\gamma$-photons and a plane-wave description of the laser field. In the 1990s, the nonlinear Breit-Wheeler process has been observed experimentally in a few-photon regime ($n\sim 5$) at $\xi\lesssim 1$, where the process rate was found to follow a power law of the form $R\sim\xi^{2n}$ [11].
@@
A promising possibility to generate high-energy $\gamma$-rays is offered by bremsstrahlung of highly relativistic electrons from high-$Z$ targets, as proposed by Reiss in the 1970s [5], and recently further investigated by several groups [35, 36, 32, 37, 33, 34, 38, 39]. In this situation, the $\gamma$-photons are not monoenergetic but span a broad range of energies. This route is followed in the experimental designs for nonlinear Breit-Wheeler pair creation at CALA [13] and DESY [14]. If successful, they could be considered as ’discovery experiments’ for the highly nonlinear regime of the strong field Breit-Wheeler process, which will be followed afterwards by ’precision experiments’ that aim at covering large regions of the parameter space with enhanced statistical accuracy. A detailed theoretical analysis of the experiment planned at CALA has recently been carried out [33, 34], including laser focusing effects and the broad frequency spectrum of bremsstrahlung. This experiment is going to operate in the $\xi\gg 1$ regime with $\kappa\sim 1$, exploiting laser wakefield accelerated incident electrons of about 2.5 GeV energy to generate the bremsstrahlung in a tungsten target [13].
@@
In the present paper, we extend the previous analysis [33] to the overcritical regime of large $\kappa$ values up to $\kappa\approx 30$. It could, for instance, be realized in the future at the Extreme Light Infrastructure (ELI) [15] or CoReLS [40] facilities. Our main goal is to reveal characteristic differences between the intermediate ($\kappa\approx 1$) and overcritical ($\kappa\gg 1$) regimes of the nonlinear strong-field Breit-Wheeler process with focused laser pulses and $\gamma$-photons from bremsstrahlung. [^1] The $\kappa$ values considered here remain far below the domain where perturbation theory with respect to the quantized radiation field is expected to break down, i.e. $\kappa\ll\alpha^{-3/2}\approx 1.6\times 10^{3}$, with $\alpha$ referring to the fine-structure constant [8, 9, 10, 41]. We will show that the overcritical field regime requires an adjusted theoretical approach and exhibits qualitatively distinct properties in terms of its scaling behaviors. This affects, in particular, its dependencies on the laser pulse duration and focusing as well as the mainly contributing spectral range of $\gamma$-photons. Our outcomes can be relevant for the design of future experiments.
@@
It is worth noting that the overcritical regime of strong-field pair production can also be probed when an incident high-energy electron beam interacts directly with an ultrastrong laser pulse [12, 42, 43]. In this case, however, the electron beam can be strongly distorted by radiation reaction effects before reaching the focal region of highest field strength. In this regard, the use of a target foil that first converts the incident electron energy in a controled way into $\gamma$-photons, which afterwards penetrate into the laser focal region, offers certain advantages. Figure 1 shows a scheme of the envisaged setup.
@@
Figure 1: Scheme of the experimental setup for entering the overcritical regime of the nonlinear Breit-Wheeler pair production process from the collision of bremsstrahlung $\gamma$-photons and a high-intensity laser pulse. The collision of a monoenergetic electron beam with a narrow high-$Z$ target will result in the bremsstrahlung burst utilized in the setup. A magnet after the target will deflect residual electrons away from the site where the photon-photon collision occurs.
@@
Our study is presented as follows: In Sec. II we first briefly review the theoretical description of Breit-Wheeler pair creation in collisions of bremsstrahlung $\gamma$-rays with a focused high-intensity laser pulse from Ref. [33] and, afterwards, extend it to the overcritical regime. Next, in Sec. III, our numerical results obtained for the overcritical regime $\kappa>1$ are discussed and put in contrast to the properties of the pair creation process at $\kappa\approx 1$. Finally, the conclusions are drawn in the last section. Throughout the manuscript, the Lorentz-Heaviside units [$c=\hbar=\epsilon_{0}=1$] and the metric tensor with the signature $\mathrm{diag}(g^{\mu\nu})=(1,-1,-1,-1)$ are used.
@@ II Theoretical treatment
Equation (1) forms the basis for two different theoretical approaches to strong-field Breit-Wheeler pair production that will be outlined in Secs. II.2 and II.3 below. They are used to calculate the first generation of pairs, which are created directly from the incident $\gamma$-photons. An extension of the approach, which allows us to include also the second generation of pairs from a shower-like process, will be described in Sec. II.4. Beforehand, we shall specify the colliding high-intensity focused laser pulse and high-energy bremsstrahlung $\gamma$-rays.
II.1 Descriptions of the focused laser pulse and bremsstrahlung spectrum
@@
where $\tau$ stands for the pulse duration (at FWHM from the intensity), $\rho^{2}=x^{2}+y^{2}$, $\zeta=z/z_{R}$ with the Rayleigh length $z_{R}=\pi w_{0}^{2}/\lambda$ and the beam waist radius at the focus $w_{0}$, whereas $w(z)=w_{0}\sqrt{1+\zeta(z)^{2}}$. Moreover, $E_{x}=B_{y}$ holds and the field phase in Eq. (2) reads
@@
for a counterpropagating geometry, in which case, the interacting $\gamma$-quantum is characterized by a wave vector $k^{\prime}=\omega^{\prime}(1,0,0,-1)$.
In order to take into account bremsstrahlung, the corresponding photon distribution has to be considered. For high energies $E_{0}$ of incident bremsstrahlung electrons—which is the case in the present paper as we assume GeV energies—the collimation angle of the photons emitted by a single electron can be approximated as an inverse Lorentz factor $\theta_{\gamma}\approx 1/\gamma_{e}=m/E_{0}\sim\mathcal{O}(1)$ mrad. Hence, the vast majority of the produced bremsstrahlung photons is emitted in the direction of propagation of incident electrons. Consequently, the photons energy spectrum can be approximated by [45, 46]
@@
with the normalized photon energy $f=\omega^{\prime}/E_{0}$ and the normalized target thickness $\ell=L_{\mathrm{T}}/L_{\mathrm{rad}}$, which relates the target thickness $L_{\mathrm{T}}$ to the radiation length of the material $L_{\rm rad}$. Note that the formula above is valid within the complete screening approximation and for materials with high $Z$-numbers, as well as for very thin targets with $\ell\ll 1$. Its divergence for $f\to 0$ is not harmful in the present context, because the low-energy part of the spectrum below $f\lesssim 0.01$ will practically not contribute to the pair creation in the considered parameter regime (see Sec. III). Besides, its non-zero value at the spectral endpoint $f=1$ does not distort the calculated pair yields either, because the real bremsstrahlung spectrum would decline to zero very steeply at this point (see Ref. [33] for more details, in particular Fig. 6 therein).
Hereafter, a target made up of tungsten is considered with $L_{\rm rad}=3.5$ mm and $L_{T}=50\>\rm\mu m$, corresponding to $\ell=0.015$ [13]. In Fig. 2 one sees the bremsstrahlung spectrum per radiating incident electron as given in Eq. (5) for the chosen target thickness as a function of $f$. We shall assume throughout that the incident electron bunch contains a total (absolute) charge of $Q_{e}=10$ pC and that $\delta=1$% of the electrons will emit bremsstrahlung, which represents a reasonable fraction for the chosen target thickness [13]. Accordingly, the total number of radiating electrons is $Q_{e}\delta/|e|\approx 6\times 10^{5}$.
@@
where $\theta_{\mathrm{rms}}=(\theta_{e^{-}}^{2}+\theta_{\gamma}^{2})^{1/2}$ is the root-mean-square value of the electron-beam divergence and the angular spread of emitted bremsstrahlung for a single electron. Here, it is assumed that the electrons are accelerated via laser wakefield acceleration by utilizing a portion of the same laser seed from which the tightly focused laser pulse for the pair production is generated. Moreover, for the longitudinal extension of the electron beam a conservative value of $2z_{R}$ is supposed (i.e. twice the Rayleigh length of the latter pulse). For further details, we refer to Sec. II.1 in [33]. In the further course of our consideration the electron beam spreading parameters are set to $L_{e^{-}}=10$ cm and $\theta_{e^{-}}=0.5$ mrad (see [13]). The distance to the interaction zone travelled by the $\gamma$-photons is taken as $L=0.5$ m. Accordingly, when the incident electron energy is $E_{0}=2.5$ GeV, the cross sectional area amounts to $A_{\gamma}\approx\pi(320\mu{\rm m})^{2}$. It slightly shrinks to $A_{\gamma}\approx\pi(300\,\mu{\rm m})^{2}$ for $E_{0}=10$ GeV, because the beam of $\gamma$-photons is less divergent then.
@@ II.2 Pair production at moderate field strengths: Stable $\gamma$-beam model
Since the transverse extension of the bremsstrahlung $\gamma$-beam is much larger than the transverse size of the laser pulse, the latter determines their overlap region and, thus, the upper boundary of the $\rho$ coordinate; in the calculations, the value $\rho_{\rm max}=3w_{0}$ has been taken. The boundaries of the $z$-integration take the spatial overlap between laser pulse and bremsstrahlung beam into account as well (see also Fig. 7 below). In order to include the energy distribution of the bremsstrahlung, an incoherent average of Eq. (8) over the spectrum (5) must be taken. Thus, the number of created Breit-Wheeler pairs per radiating electron in our S-model is given by
@@ II.3 Pair production at very high field strengths: Decaying $\gamma$-beam model
The method described in Sec. II.2 is only applicable when the quantum nonlinearity parameter $\kappa$ and the laser pulse duration $\tau$ are not too large. Otherwise this method will overestimate the pair production yield, because it disregards the attenuation of the $\gamma$-rays that arises from their decay into electron-positron pairs. In the following we present another calculational method that is able to describe the nonlinear Breit-Wheeler process at $\kappa$ values substantially exceeding unity. To this end, we adjust the treatment put forward in Ref. [32] and extend it to incorporate $\gamma$-rays from bremsstrahlung with a broad frequency spectrum.
@@
which coincides with the probability for producing a pair. In the exponent of Eq. (12), the pair production rate $R(\kappa)$ enters, which is evaluated along the trajectory of the $\gamma$-photon. We note that $R(\kappa)$ is valid up to the first order in $\alpha$ [see Eq. (1)]. Higher-order processes such as $\omega^{\prime}+n\omega\to e^{+}e^{-}+\omega^{\prime\prime}$, where the pair production is accompanied by the emission of a $\gamma$-photon $\omega^{\prime\prime}$, are not included here. We will come back to this point in Sec. II.4 below.
Clearly, when the time-integrated rate in Eq. (12) is much smaller than unity, the exponential can be expanded and one obtains
@@
Afterwards, we obtain the total number of produced pairs per radiating electron in the D-model by weighting with the bremsstrahlung spectrum, according to
@@
For our numerical calculations we cut off the integration domain in Eq. (14) at a sufficiently large upper boundary $b_{\rm max}=3w_{0}$, that corresponds to a cross sectional area of $A_{\rm max}=\pi(3w_{0})^{2}$. Then we distribute the associated portion of $A_{\rm max}/A_{\gamma}$ bremsstrahlung $\gamma$-photons uniformly over the area $A_{\rm max}$ and calculate the contribution of each trajectory to the pair yield. It is worth mentioning that the maximal value of $\kappa$ [see Eq. (4)] is experienced by those $\gamma$-photons which pass through the center of the laser focus at $b\approx 0$ and have an energy close to $E_{0}$.
II.4 Extension of the D-model:
Second generation of created pairs
Our description of the production process in Secs. II.2 and II.3 solely accounts for an initial pair creation event, i.e. for the first generation of created pairs. Since the particles are generated with high energies and, directly after their creation, are still subject to the strong laser field, they can in principle induce further pair production events. In particular, a created high-energy electron (or positron) counterpropagating a high-intensity laser pulse can emit a $\gamma$-photon through Compton scattering, which afterwards induces another Breit-Wheeler pair creation event. Such pair production showers (or cascades) have been studied intensively in recent years [47, 48, 49, 50, 51].
By a suitable amendment of our treatment, we shall account—in an approximate way—for the possibility of subsequent pair creation events which are induced by the first generation of pairs. Previous calculations of nonlinear Breit-Wheeler pair production by a monoenergetic $\gamma$-beam colliding with an intense laser pulse [32] have shown that the second generation of pairs starts to become relevant for values $\kappa\gtrsim 4$–5 of the quantum nonlinearity parameter. A subsequent study [51] has revealed that even higher generations of pairs would contribute for $\kappa\gtrsim 16$–20. The latter study assumed linearly polarized, plane-wave laser pulses with sin2-shaped intensity profile. The yield of created pairs was obtained by numerically solving a coupled system of integro-differential equations for the distributions functions of electrons, positrons and $\gamma$-photons (see Sec. II.C in [51]). The corresponding numerical code has been made available [52]. This way, the numbers $N_{1}$ and $N_{2}$ of pairs created in the first and second generation, respectively, were computed, from which enhancement factors $(N_{1}+N_{2})/N_{1}$ can be obtained. They quantify the amplification of the pair yields due to the second generation over the sole contribution of the first generation.
We have combined the framework developed in [51] with our treatment to estimate the additional contribution to the pair yield when the second generation is included. In a setup with incident electron energy $E_{0}$ and laser intensity parameter $\xi$, we determine the enhancement factors $(N_{1}+N_{2})/N_{1}$ for each value of $f$ and $b$ by running the code [52] multiple times. Calculating the enhancement for fixed $\gamma$-photon energy $\omega^{\prime}=fE_{0}$ and impact parameter $b$ (this way fixing the maximum field strength $\mathcal{E}_{0}(b)=\mathcal{E}_{0}{\rm e}^{-(b/w_{0})^{2}}$ experienced by the $\gamma$-photon) is necessary since the analysis in [51] assumes a monoenergetic $\gamma$-beam and unfocused laser pulse. Accordingly, by running the code separately for each pair of parameter values $\omega^{\prime}$ and $\xi(b)=|e|\mathcal{E}_{0}(b)/(m\omega)$, we scan over the bremsstrahlung spectrum and the transverse laser profile. The resulting enhancement factors are inserted into the two-dimensional integral over $f$ and $b$ [see Eqs. (14) and (15)], in order to obtain the pair yield including the contribution from the second generation.
Our corresponding results will be shown in Sec. III.2 where pair creation in collisions of bremsstrahlung with laser pulses at $\kappa\gg 1$ is considered.